# Detecting ionized bubbles around luminous sources during the reionization era using HI 21-cm signal

### v2

@@ 5.5 Constraining the neutral hydrogen fraction
To calculate $C$, we simulate an entirely neutral and uniform HI 21-cm signal around a spherical ionized bubble of radius 23 Mpc, i.e., $x_{\rm HI}=1$ outside the ionized bubble. It allows us to estimate $C$ using eq. 5.1. Now, after adding the foreground component only with the HI 21-cm signal, we follow the entire analysis and calculate the SNR as a function of filter size without adding system noise from the radio interferometers. The SNR, in this case, peaks at 23 Mpc and is supposed to be equal to $C$ (following eq. 5.1) as the $x_{\rm HI}=1$ outside the ionized bubble in this case. We use this value of $C$ to estimate the mean $x_{\rm HI}$ and the $1\sigma$ uncertainty associated with it. The estimated neutral fraction is $x_{\rm HI}=0.93\pm 0.15$ at $z=7.1$ for $2048$ hrs of observations with uGMRT. It is very close to the actual $x_{\rm HI}$ in the simulation, which is $0.88$. The estimated $1\sigma$ upper limit of $x_{\rm HI}$ exceeds unity which is, of course, unrealistic. The SKA1-Low places much tighter constraints, which is $x_{\rm HI}=0.92\pm 0.055$ using only $96$ hrs of observations for the same scenario. Similarly we also studied this for the third case, where the neutral hydrogen fraction is lower in the simulation, discussed in subsection 5.3. We find the estimated neutral hydrogen fraction to be $\bm{x_{\rm HI}=0.49\pm 0.09}$ at $\bm{z=7.1}$ for $\bm{5120}$ hours of observations with the uGMRT, which is very close to the actual neutral hydrogen fraction $\bm{x_{\rm HI}=0.52}$.
@@ 6 Summary & Discussion
First, we simulated the HI 21-cm signal around bright QSOs/galaxies matching the above cases and foreground contaminants in a 3D cube. Subsequently, we have generated visibilities corresponding to the HI 21-cm signal and foregrounds at baselines similar to the uGMRT and SKA1-Low experiments and added the system noise with the visibilities corresponding to the HI signal and foreground contaminants. We then employ a foreground mitigation technique based on a polynomial fitting algorithm to remove them from the total visibilities. The residual visibilities match well with the combination of the input HI signal and system noise. However, it is heavily dominated by the system noise. To minimize the effects of system noise in the estimated quantity, we employed a technique based on the matched filter algorithm. This technique is applicable in our case as the functional form of the desired HI 21-cm signal around bright sources is known. We find that the expected signal-to-noise ratio for the proposed estimator for the uGMRT is $\bm{\sim 5}$ for both $\bm{\sim 2000}$ hours of observations at redshift $\bm{7.1}$ and $\bm{\sim 3000}$ hours at redshift $\bm{8.3}$, when the mean HI fraction outside the targeted bubble is $\bm{\sim 0.9}$. However, the required observing time increases to $\bm{\sim 5120}$ hrs for the uGMRT at redshift $\bm{7.1}$ to achieve a similar SNR when the mean HI fraction is lower at $\bm{\sim 0.5}$. The SKA1-Low will have a much higher SNR ($\sim 8-10$) with just $\sim 100$ hrs of observations at same redshifts $7.1$ and $8.3$, when the mean HI fraction outside the targeted bubble is $\bm{\sim 0.9}$. However, the required observing time increases to $\bm{\sim 200}$ hrs for the SKA1-Low at redshift $\bm{7.1}$ to achieve a similar SNR when the mean HI fraction is lower at $\bm{\sim 0.5}$.
Apart from detecting the HI 21-cm signal, the technique is also able to measure the size of the ionized regions and HI neutral fraction. For the uGMRT $\sim 2000$ hours of observations at redshift $7.1$, we find that the estimated bubble size ranges between $22-26$ Mpc for $100$ independent noise realizations when the actual bubble size is $23.5$ Mpc. Similar results are obtained for $\bm{\sim 5120}$ hours of observations with the uGMRT at redshift $\bm{7.1}$ with a mean HI fraction of $\bm{\sim 0.5}$ and for $\bm{\sim 3000}$ hours at redshift $\bm{8.3}$ with a mean HI fraction of $\bm{\sim 0.9}$. The SKA1-Low should measure the bubble size more accurately, with the uncertainty in bubble size estimation being less than $1$ Mpc at both mentioned redshifts.
@@
Here, we focused on the targeted detection of ionized bubbles around sources previously detected by other telescopes. Several such sources have been observed during the reionization epoch. Additionally, the matched filter technique can detect ionized bubbles even without prior source detection. However, this unguided detection requires exploring a vast parameter space, which is computationally expensive. The underlying bubble may not be spherical, as assumed in our study. However, as long as the bubble is relatively isolated, the method remains applicable, though the SNR might be reduced. Foregrounds are assumed to be spectrally smooth here, which may not be accurate in actual observations. A more sophisticated method for foreground mitigation may be needed in case foregrounds behave differently. We would also like to note that this study does not account for potential errors arising from improper calibration [64], ionospheric effects, or radio-frequency interference (RFI). These practical issues will be addressed in future works.

### v3

@@ 5.5 Constraining the neutral hydrogen fraction
To calculate $C$, we simulate an entirely neutral and uniform HI 21-cm signal around a spherical ionized bubble of radius 23 Mpc, i.e., $x_{\rm HI}=1$ outside the ionized bubble. It allows us to estimate $C$ using eq. 5.1. Now, after adding the foreground component only with the HI 21-cm signal, we follow the entire analysis and calculate the SNR as a function of filter size without adding system noise from the radio interferometers. The SNR, in this case, peaks at 23 Mpc and is supposed to be equal to $C$ (following eq. 5.1) as the $x_{\rm HI}=1$ outside the ionized bubble in this case. We use this value of $C$ to estimate the mean $x_{\rm HI}$ and the $1\sigma$ uncertainty associated with it. The estimated neutral fraction is $x_{\rm HI}=0.93\pm 0.15$ at $z=7.1$ for $2048$ hrs of observations with uGMRT. It is very close to the actual $x_{\rm HI}$ in the simulation, which is $0.88$. The estimated $1\sigma$ upper limit of $x_{\rm HI}$ exceeds unity which is, of course, unrealistic. The SKA1-Low places much tighter constraints, which is $x_{\rm HI}=0.92\pm 0.055$ using only $96$ hrs of observations for the same scenario. Similarly we also studied this for the third case, where the neutral hydrogen fraction is lower in the simulation, discussed in subsection 5.3. We find the estimated neutral hydrogen fraction to be ${x_{\rm HI}=0.49\pm 0.09}$ at $z=7.1$ for $5120$ hours of observations with the uGMRT, which is very close to the actual neutral hydrogen fraction ${x_{\rm HI}=0.52}$.
@@ 6 Summary & Discussion
First, we simulated the HI 21-cm signal around bright QSOs/galaxies matching the above cases and foreground contaminants in a 3D cube. Subsequently, we have generated visibilities corresponding to the HI 21-cm signal and foregrounds at baselines similar to the uGMRT and SKA1-Low experiments and added the system noise with the visibilities corresponding to the HI signal and foreground contaminants. We then employ a foreground mitigation technique based on a polynomial fitting algorithm to remove them from the total visibilities. The residual visibilities match well with the combination of the input HI signal and system noise. However, it is heavily dominated by the system noise. To minimize the effects of system noise in the estimated quantity, we employed a technique based on the matched filter algorithm. This technique is applicable in our case as the functional form of the desired HI 21-cm signal around bright sources is known. We find that the expected signal-to-noise ratio for the proposed estimator for the uGMRT is ${\sim 5}$ for both ${\sim 2000}$ hours of observations at redshift $7.1$ and ${\sim 3000}$ hours at redshift $8.3$, when the mean HI fraction outside the targeted bubble is ${\sim 0.9}$. However, the required observing time increases to ${\sim 5120}$ hrs for the uGMRT at redshift $7.1$ to achieve a similar SNR when the mean HI fraction is lower at ${\sim 0.5}$. The SKA1-Low will have a much higher SNR ($\sim 8-10$) with just $\sim 100$ hrs of observations at same redshifts $7.1$ and $8.3$, when the mean HI fraction outside the targeted bubble is ${\sim 0.9}$. However, the required observing time increases to ${\sim 200}$ hrs for the SKA1-Low at redshift $7.1$ to achieve a similar SNR when the mean HI fraction is lower at ${\sim 0.5}$.
Apart from detecting the HI 21-cm signal, the technique is also able to measure the size of the ionized regions and HI neutral fraction. For the uGMRT $\sim 2000$ hours of observations at redshift $7.1$, we find that the estimated bubble size ranges between $22-26$ Mpc for $100$ independent noise realizations when the actual bubble size is $23.5$ Mpc. Similar results are obtained for ${\sim 5120}$ hours of observations with the uGMRT at redshift $7.1$ with a mean HI fraction of $\sim 0.5$ and for ${\sim 3000}$ hours at redshift $8.3$ with a mean HI fraction of ${\sim 0.9}$. The SKA1-Low should measure the bubble size more accurately, with the uncertainty in bubble size estimation being less than $1$ Mpc at both mentioned redshifts.
@@
Here, we focused on the targeted detection of ionized bubbles around sources previously detected by other telescopes. Several such sources have been observed during the reionization epoch. Additionally, the matched filter technique can detect ionized bubbles even without prior source detection. However, this unguided detection requires exploring a vast parameter space, which is computationally expensive. The underlying bubble may not be spherical, as assumed in our study. However, as long as the bubble is relatively isolated, the method remains applicable, though the SNR might be reduced. Furthermore, we note that our simulations do not account for the contributions of low-mass halos to reionization. This omission likely results in an underestimation of the diffuse and complex structure of the ionized regions. Accurately incorporating these contributions could, to some extent, degrade the predicted signal-to-noise ratio. Foregrounds are assumed to be spectrally smooth here, which may not be accurate in actual observations. A more sophisticated method for foreground mitigation may be needed in case foregrounds behave differently. We would also like to note that this study does not account for potential errors arising from improper calibration [64], ionospheric effects, or radio-frequency interference (RFI). These practical issues will be addressed in future works.